OpenStudy (anonymous):

On the three consecutive passes, a football team gains 10 yards, losses 29 yards, and gains 35 yards. What number represents the total net yardage?

OpenStudy (anonymous):

10-29=-19 -19+35=16 :D easy peasy

OpenStudy (anonymous):

Thank you.